OpenStudy (anonymous):

find the exact values of the six trigonometric functions of t. (-1/5, 2 sqrt 6/5)

OpenStudy (anonymous):

this is a point on the unit circle since \[(\frac{1}{5})^2+(\frac{2\sqrt{6}}{5})^2=1\]

OpenStudy (anonymous):

so you can read them off sine is the second coordinate cosine is the first tangent is second over first cosecant is the reciprocal of the second secant is the reciprocal of the first cotangent is the reciprocal of tangent or first over second

OpenStudy (anonymous):

I'm a little confused still....would you mind explaining again?